# Aligning Visual Contrastive learning models via Preference Optimization

### v1

@@ 3.4 Linear Transformations and Adaptations
In this formulation, the unitary matrix $V$ selects the important directions, and the diagonal matrix $\Sigma=\text{diag}(\sigma_{1},\dots,\sigma_{n})$ determines the degree of attenuation or amplification along each direction $v_{i}$. Specifically, each $\sigma_{i}$ determines whether the corresponding direction $v_{i}$ is strengthened ($\sigma_{i}>1$) or weakened ($\sigma_{i}<1$). Since our training process keeps the model close to the reference, each $\sigma_{i}$ remains relatively close to 1, and the overall matrix $W^{T}W$ remains close to the identity matrix.
This observation allows for post-training adjustments, where we can strengthen or weaken each $v_{i}$ as desired. One method for achieving this is by directly tuning each singular value $\sigma_{i}$ using either linear interpolation or matrix powers, as follows:
@@ 4 Experiments and Results
In this section, we experimentally justify our method capabilities, using the CLIP model as our contrastive learning model. An explanation of the datasets we used can be found in section E.
Setup: For all experiments, we used 8 A100 GPUs, each with 40GB of memory. For the typographic attack experiments, hyperparameters such as $\beta$ and $\lambda$ were selected based on the ablation studies in the appendix. Specifically, we used $\beta=\lambda=1$ for DPO, $\beta=\lambda=0.01$ for IPO, and $\beta=1.5$, $\lambda=0.01$ for KTO. In disentangling gender bias, we also set $\beta=\lambda=1$. All models were trained for three epochs with a batch size of 512. The learning rate was set to $2\times 10^{-5}$, and we employed the Adamax optimizer with a linear warmup and a cosine scheduler, setting the warmup ratio to 0.1. We set the random seed to 0 for all experiments to ensure reproducibility.
@@ 4.1 Typographic Robustness
This section evaluates our method using the typographic attack mitigation task. We set up the problem as described, using the template "an image of a <class $i$>" (or a more dataset-specific template for non-generic images), designating the typographic class as the undesired outcome $y_{l}$ and the original caption as the desired outcome $y_{w}$.
Baselines: We evaluate our method against the following baselines: pre-trained CLIP (Radford et al., 2021), PAINT (Ilharco et al., 2022), and Defense-prefix (Azuma & Matsui, 2023b), with results collected from (Azuma & Matsui, 2023b) for comparability. We use ImageNet-100, a 100-class subset of ImageNet (Russakovsky et al., 2015) to train the model. Typographic attack images are generated by adding misleading text labels to the original images, using open-source implementations from (Azuma & Matsui, 2023b).
Table 1: Classification accuracy results on the datasets: O (Original dataset) and T (Typographic dataset). Note that other methods’ analyses and experimental results on the ImageNet dataset were conducted on a subset called ImageNetV2 validation, which was not publicly available when these results were derived. We report the results on different datasets.
Method $\text{{ImageNet}}^{*}$ Caltech101 OxfordPets StanfordCars Flowers102 FGVCAircraft DTD SUN397 EuroSAT
O T O T O T O T O T O T O T O T O T
CLIP 62.02 39.10 88.64 63.97 87.35 58.95 58.72 21.02 66.32 31.32 18.99 10.83 44.57 25.53 61.74 34.02 42.98 4.86
Materzynska+ 54.38 44.91 80.53 74.73 75.01 63.61 40.33 15.79 51.86 34.95 13.23 8.28 36.28 33.03 51.06 39.52 37.32 16.22
PAINT 61.82 55.90 88.48 83.57 85.23 76.53 55.30 33.44 64.73 54.92 17.73 14.46 42.61 36.60 61.69 53.62 38.20 17.31
Defense-Prefix 62.48 49.83 89.28 79.54 87.22 72.86 57.47 28.64 63.82 44.12 19.26 14.49 40.64 31.60 61.41 43.50 43.85 9.85
Ours (DPO) 53.80 49.18 87.50 85.43 85.25 79.72 56.03 34.33 56.60 55.70 16.21 13.87 39.36 38.48 61.02 56.34 49.33 28.32
Ours (IPO) 47.61 42.75 85.73 83.78 85.32 80.44 53.67 35.02 54.50 52.80 17.97 15.86 40.53 39.94 61.91 58.05 46.12 43.23
Ours (KTO) 53.17 50.35 87.67 86.02 85.41 81.02 57.76 37.04 59.10 58.00 17.27 15.59 40.74 40.33 62.52 59.01 46.26 36.94
Difference $\downarrow$8.68 $\downarrow$5.55 $\downarrow$1.61 $\uparrow$2.45 $\downarrow$1.94 $\uparrow$4.49 $\downarrow$0.96 $\uparrow$3.60 $\downarrow$5.63 $\uparrow$3.08 $\downarrow$1.29 $\uparrow$1.37 $\downarrow$1.87 $\uparrow$3.73 $\uparrow$0.17 $\uparrow$5.39 $\uparrow$5.84 $\uparrow$25.92
Our results, as seen in table 1, indicate that our method can effectively prevent typographic attacks with little loss of prior knowledge. In addition, we provided a qualitative analysis in Appendix C and D.
Table 2: Classification accuracy results on real typographic datasets.
@@ 4.1.1 Control between Optical Character Recognition and Object detection
In this experiment, we apply linear adaptations and retrain the network, then evaluate the training results. We then interpolate between models using previously discussed linear transformation methods, varying the transformation scaling $t$, defined in 15 from $-2.0$ to $1.2$ to assess performance. By doing this, we measure the effectiveness of our controlling scheme between the Optical Character Recognition (OCR) and Object Detection tasks.
Figure 2: (Left) Accuracy on typographic samples and percentage of typographic label predictions versus transformation scaling factor $t$. As $t$ increases, the model favors object labels over typographic labels while maintaining accuracy. (Right) Frontier of a DPO fine-tuned model, showing OCR accuracy vs. object detection accuracy across models with varying $t$.
Our results in Figure 2 demonstrate that we can confidently control the trade-off between OCR and object detection tasks without any additional training and with minimal impact on overall accuracy. This method allows us to choose which task to prioritize flexibly, and we can visualize the performance frontier, showing that improvements in OCR accuracy come at the cost of object detection accuracy and vice versa. The ability of this linear layer to manage the trade-off suggests that the concept is linearly separable within the CLIP embedding space.
@@ 4.2 Disentangling Gender Understanding
In this setup, we demonstrate how our approach can completely reverse a contrastive learning model’s understanding of a concept, such as gender, without significantly altering other aspects of the model’s performance.
(I)(II)Probabilities Across Models
@@
Transformation Scaling (t)
Figure 3: (Left) Image input. (Center) Comparison of model predictions before and after applying our gender-flipping method, showing changes in the predicted gender probabilities. (Right) Performance across transformation scaling $t$: As $t$ increases from 0 to 1, gender-specific predictions are reversed. The intersection at $t^{*}$ marks the point where gender information is effectively neutralized, leading to balanced male and female predictions. Task detection accuracy remains stable across all models. (I) and (II) are the corresponding captions: (I) The prediction that the man in the photo is working. (II) The prediction is that the woman in the photo is working.
[…]
As the regulatory term, we use the same dataset of images. However, we employ androgynous target labels such "The person in <activity>" to ensure that our model does not forget any information on detecting the correct activity. We repeat our learning process using only the linear layer and then visualize the effect of varying the transformation scaling $t$ between 0 and 1. Our results confirm that we have a high degree of control over CLIP’s understanding of gender, as demonstrated in Figure 3.
To further evaluate our results, we apply the modified CLIP models to an image retrieval task, distinct from the original text retrieval task used for training. We assess how this impacts the retrieved images for specific text prompts by computing the similarity scores between the prompts and images as $s(\text{image},\text{text})=\mathcal{I}_{\theta}(\text{image})^{T}\mathcal{T}_{\theta}(\text{text})$. Using the dataset from (Zhang et al., 2022b), we issue prompts describing individuals in various occupations to analyze the model’s behavior and showcase an instance in Figure 4.
Figure 4: Retrieved images for caption ”an image of a police”, Top: Reversed(6W,4M), Middle: Original(2W, 8M), Bottom: Neutralized(5W, 5M) being the model at $t=t^{*}$
@@ 4.3 Preserving Pretrained Knowledge via Preference Optimization
In all of the aforementioned preference optimization methods, an implicit or explicit constraint exists to keep the model as close as possible to the reference model. Specifically, including KL-divergence regularization facilitates adjustments without deviating significantly from the reference model $\pi_{\text{ref}}$. To support this claim, Figure 5-(a) shows the KL-divergence between the output distributions of the model and the reference model on the original dataset, measured as $\mathbb{E}_{x\sim\mathcal{D}_{\text{reg}}}\left[D_{\mathrm{KL}}(\pi_{\theta}(\cdot|x)||\pi_{\mathrm{ref}}(\cdot|x))\right]$, after fine-tuning for the typographic robustness task using different optimization methods.
Compared to Cross-Entropy optimization, our proposed methods result in lower divergence, demonstrating better retention of the model’s pre-trained knowledge. Additionally, Figure 5-(b) illustrates that we can control this deviation (i.e., retention of pre-trained knowledge) during fine-tuning by adjusting the $\beta$ hyperparameter, consistent with Equation 3. All models in (a) and (b) were trained on the Imagenet100 and FOOD101 datasets, respectively, and the models in (b) are optimized using the IPO method.
Refer to caption Refer to caption
(a) (b)
Figure 5: (a) KL-divergence comparison of our methods versus Cross-Entropy optimization. (b) Effect of the hyperparameter $\beta$ on divergence between the model and reference model.
@@ 5 Conclusion
This paper introduced a novel framework for improving contrastive learning models through Preference Optimization (PO). Our approach systematically aligns model behavior with desired preferences, enhancing robustness against typographic attacks and effectively mitigating biases, such as gender bias, in vision-language models like CLIP. By extending techniques such as Direct Preference Optimization (DPO) and Identity Preference Optimization (IPO), we showed that preference-based optimization can fine-tune models to meet specific behavioral requirements while maintaining their pre-trained knowledge. Our experiments demonstrated improved performance over standard contrastive learning methods in adversarial robustness and fairness, particularly in handling sensitive attributes like gender and race.
This work could open new pathways for applying preference-based optimization to non-generative models. It highlights its potential for tasks requiring fairness, alignment with human values, and robustness to adversarial attacks. Future work could explore further extensions of PO to other domains and modalities and investigate more fine-grained control mechanisms for tuning model behaviors post-training.
@@ Appendix A Theoretical Foundations of RLHF
RLHF extends the traditional reinforcement learning paradigm by incorporating human preferences into the learning process (Ouyang et al., 2022a). This feedback can take the form of direct preference comparisons between different actions or sequences of actions, or it can involve explicit corrections when the agent’s behavior is undesirable. Human judgments are converted into a reward signal that helps the agent refine its policy. RLHF has gained attention in areas like natural language processing and fine-tuning large language models, where specifying a clear reward function is difficult and aligning the agent’s behavior with human values or intentions is critical. This method offers a more robust approach to training AI systems in tasks that require subjective judgment or nuanced decision-making.
@@
where $D_{\mathrm{KL}}(\pi_{\theta}^{\text{RL}}(y|x)\|\pi_{\theta}^{\text{SFT}}(y|x))$ represents the KL divergence between the RL policy and the SFT policy. The coefficient $\beta$ controls the trade-off between encouraging exploration and maintaining consistency with the supervised model. L divergence serves two purposes: it acts as an entropy bonus, encourages exploration, and ensures that the RL policy stays within the supervised fine-tuned model.
@@ A.1 Deriving Preference Optimization as a Supervised Learning Framework
Direct Preference Optimization (DPO). DPO (Rafailov et al., 2023) leverages offline preference data to directly optimize the policy without relying on reinforcement learning-based methods like PPO. DPO demonstrates that the optimal solution to Eq. 19, $\pi_{\theta}^{*}$, satisfies the following equation:
@@
The preference probability $P_{\theta}$ is derived from pairwise comparisons using the Bradley-Terry model. Substituting Equation 20 into Equation 18 yields the following loss function for DPO:
@@
In Equation 21, $\sigma$ denotes the sigmoid function, and $\mathcal{D}$ represents the dataset consisting of pairwise preferences. Each preference triplet $(x,y_{w},y_{l})$ includes a prompt $x$, a preferred response $y_{w}$, and a less preferred response $y_{l}$. This formulation allows DPO to optimize directly on pairwise preference data, avoiding the instability of RL methods while maintaining mathematical equivalence with traditional RLHF.
Identity Preference Optimization (IPO). As the original DPO loss function has shown limitations in practice, such as overfitting to the preference dataset (Azar et al., 2023a), IPO was proposed as an extension. IPO introduces a regularization term to the DPO loss to mitigate overfitting by controlling the gap between the log-likelihood ratios of the preferred and dispreferred outputs for both the model and the reference model. The IPO loss function is defined as:
@@
Kahneman-Tversky Optimization (KTO). In language modeling, collecting large-scale preference datasets can be difficult, but it is relatively easier to determine whether an output is desirable or undesirable. This challenge has inspired KTO (Kawin Ethayarajh et al., 2024), a method that bypasses the need for detailed preference data by using a binary signal to guide optimization. KTO is based on prospect theory (Tversky, 2016), directly maximizing the utility of generations using a model of human decision-making.
@@
where $v_{\text{KTO}}(x,y;\beta)$ adjusts based on whether the output is desirable or not:
@@
Here, $z_{\text{ref}}$ represents the reference reward, and $r_{\text{KTO}}(x,y)=\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}$. The weighting function $w(y)$ differentiates between desirable and undesirable outputs, where $\lambda_{D}$ and $\lambda_{U}$ are hyper-parameters controlling the weighting of each.
KTO’s key advantage is its ability to focus on maximizing the reward for desirable outputs without inflating the KL divergence term, effectively balancing utility and regularization. This makes KTO a practical solution for real-world applications, where collecting detailed preference data is infeasible.

### v3

@@ 3.4 Linear Transformations and Adaptations
In this formulation, the unitary matrix $V$ selects the important directions, and the diagonal matrix $\Sigma=\text{diag}(\sigma_{1},\dots,\sigma_{n})$ determines the degree of attenuation or amplification along each direction $v_{i}$. Specifically, each $\sigma_{i}$ determines whether the corresponding direction $v_{i}$ is strengthened ($\sigma_{i}>1$) or weakened ($\sigma_{i}<1$). As detailed in Appendix E, our training process keeps the model close to the reference, as expected. Consequently, each $\sigma_{i}$ remains relatively close to 1, and the overall matrix $W^{T}W$ stays close to the identity matrix.
This observation enables post-training adjustments, allowing us to strengthen or weaken each $v_{i}$ as needed. One approach to achieve this is by directly tuning each singular value $\sigma_{i}$, either through linear interpolation or by applying matrix powers. In our work, we utilize matrix powers to modify the singular values via a transformation scaling parameter $t\in\mathbb{R}$, implementing the transformation $\Sigma\to\Sigma^{t}$ as follows:
@@
For large positive or negative values of $t$, we recommend normalizing the output embeddings, as done in the original CLIP model. This normalization prevents certain directions in the embedding space from becoming excessively large. While normalization still allows for the strengthening or weakening of specific directions, it ensures that the length of the vector remains bounded, instead of becoming arbitrarily large. This approach maintains the intended alignment or misalignment with certain directions without causing instability in the embedding space.
@@ 4 Experiments and Results
In this section, we experimentally justify our method capabilities, using the CLIP model as our contrastive learning model. An explanation of the datasets we used can be found in Appendix F.
Table 1: Classification accuracy on various datasets: O (Original dataset) and T (Typographic dataset). The last row highlights the differences between KTO, our best-performing proposed method, and the best-performing baseline for each dataset. Positive differences indicate improvements achieved by the proposed method.
Method Caltech101 OxfordPets StanfordCars Flowers102 FGVCAircraft DTD SUN397 EuroSAT Avg.
O T O T O T O T O T O T O T O T O T
CLIP 88.64 63.97 87.35 58.95 58.72 21.02 66.32 31.32 18.99 10.83 44.57 25.53 61.74 34.02 42.98 4.86 58.66 31.31
Materzynska+ 80.53 74.73 75.01 63.61 40.33 15.79 51.86 34.95 13.23 8.28 36.28 33.03 51.06 39.52 37.32 16.22 48.25 35.77
PAINT 88.48 83.57 85.23 76.53 55.30 33.44 64.73 54.92 17.73 14.46 42.61 36.60 61.69 53.62 38.20 17.31 56.74 46.31
Defense-Prefix 89.28 79.54 87.22 72.86 57.47 28.64 63.82 44.12 19.26 14.49 40.64 31.60 61.41 43.50 43.85 9.85 57.87 40.58
Ours (DPO) 87.50 85.43 85.25 79.72 56.03 34.33 56.60 55.70 16.21 13.87 39.36 38.48 61.02 56.34 49.33 28.32 56.41 49.02
Ours (IPO) 85.73 83.78 85.32 80.44 53.67 35.02 54.50 52.80 17.97 15.86 40.53 39.94 61.91 58.05 46.12 43.23 55.72 51.14
Ours (KTO) 87.67 86.02 85.41 81.02 57.76 37.04 59.10 58.00 17.27 15.59 40.74 40.33 62.52 59.01 46.26 36.94 57.09 51.74
Difference $\downarrow$1.61 $\uparrow$2.45 $\downarrow$1.81 $\uparrow$4.47 $\uparrow$0.9 $\uparrow$3.60 $\downarrow$5.63 $\uparrow$3.08 $\downarrow$1.99 $\uparrow$1.10 $\downarrow$1.87 $\uparrow$3.73 $\uparrow$0.83 $\uparrow$5.39 $\uparrow$2.41 $\uparrow$19.63 $\downarrow$0.78 $\uparrow$5.43
Setup: For all experiments, we used 8 A100 GPUs, each with 40GB of memory. For the typographic attack experiments, hyperparameters such as $\beta$ and $\lambda$ were selected based on our empirical studies described in the appendix. Specifically, we used $\beta=\lambda=1$ for DPO, $\beta=\lambda=0.01$ for IPO, and $\beta=1.5$, $\lambda=0.01$ for KTO. In disentangling gender bias, we also set $\beta=\lambda=1$. All models were trained for three epochs with a batch size of 512. The learning rate was set to $2\times 10^{-5}$, and we employed the Adamax optimizer with a linear warmup and a cosine scheduler, setting the warmup ratio to 0.1. We set the random seed to 0 for all experiments to ensure reproducibility.
@@ 4.1 Typographic Robustness
We further evaluate our method on the typographic attack mitigation task. We set up the problem as described in Section 3, using the template "an image of a <class $i$>" (or a more dataset-specific template for non-generic images), designating the typographic class as the undesired outcome $y_{l}$ and the original caption as the desired outcome $y_{w}$.
Baselines: We evaluate our method against the following baselines: pre-trained CLIP (Radford et al., 2021), (Materzynska et al., 2022), PAINT (Ilharco et al., 2022), and Defense-prefix (Azuma & Matsui, 2023), with results collected from (Azuma & Matsui, 2023) for comparability. (Materzynska et al., 2022) propose using orthogonal projections to create subspaces in CLIP’s image encoder that separate the processing of written words from visual concepts. PAINT (Patching with Interpolation) is a technique for adapting open-vocabulary models to downstream tasks, while Defense-Prefix is specifically designed to counter typographic attacks on CLIP by inserting a learned token before class names in text prompts. We use ImageNet-100, a 100-class subset of ImageNet (Russakovsky et al., 2015), to train the model. Typographic attack images are generated by adding misleading text labels to the original images using open-source implementations from (Azuma & Matsui, 2023). Further details of our baselines are provided in Appendix F.
Our results shown in Table 1 indicate that our methods effectively prevent typographic attacks with little loss of prior knowledge, with improvements of up to $19.63$ in accuracy over the best performing baseline. [^5] More analyses are shown in Appendices C, D and J.
@@ 4.1.1 Control between Optical Character Recognition and Object detection
In this experiment, we apply linear adaptations and retrain the network, then evaluate the training results. We then interpolate between models using previously discussed linear transformation methods, varying the transformation scaling $t$, defined in Eq. (16) from $-4.0$ to $4.0$ to assess performance. By doing this, we measure the effectiveness of our controlling scheme between the tasks of Optical Character Recognition (OCR) and Object Detection (OD).
Our results in Figure 2 demonstrate that we can confidently control the trade-off between OCR and OD without any additional training and with minimal impact on overall accuracy. This method allows us to choose which task to prioritize flexibly, and we can visualize the performance frontier, showing that improvements in OCR accuracy come at the cost of OD accuracy and vice versa. The ability of this linear layer to manage the trade-off suggests that the concept is linearly separable within the CLIP embedding space.
(a) Accuracy on typographic samples and percentage of typographic label predictions versus transformation scaling factor $t$. As $t$ increases, the model favors object labels over typographic labels while maintaining accuracy.
(b) Frontier of a DPO fine-tuned model, showing OCR vs. OD accuracy across with varying $t$.
Figure 2: Comparisons of optical character recognition (OCR) and object detection (OD).
@@ 4.2 Disentangling Gender Understanding
In this setup, we demonstrate how our approach can reverse a contrastive learning model’s understanding of a concept (e.g., gender) without significantly altering other aspects of its performance.
[…]
(a) Example image showing a man working.
(I)(II)
@@
(b) Model predictions before and after applying our gender-flipping method, showing changes in the predicted captions:
(I) “The man in the photo is working.“
(II) “The woman in the photo is working.“
@@
Transformation Scaling ($t^{*}$
(c) As $t$ increases from 0 to 1, gender-specific predictions are reversed. $t^{*}$ marks the point where gender information is neutralized, leading to balanced male and female predictions. Task detection accuracy remains stable across all models.
Figure 3: Analyses of the models’ understanding of gender.
To further evaluate our results, we apply the modified CLIP models to an image retrieval task, distinct from the original text retrieval task used for training. We assess how this impacts the retrieved images for specific text prompts by computing the similarity scores between the prompts and images as $s(\text{image},\text{text})=\mathcal{I}_{\theta}(\text{image})^{T}\mathcal{T}_{\theta}(\text{text})$. Using the dataset from (Zhang et al., 2022), we issue prompts describing individuals in various occupations to analyze the model’s behavior and showcase an instance in Figure 4.
Figure 4: Images retrieved for the caption ”an image of a police” with three different policies from top to bottom: reversed understanding of gender (6W, 4M), pretrained CLIP model (2W, 8M), neutralized understanding of gender (5W, 5M), i.e., $t=t^{*}$.
@@ 4.3 Preserving Pretrained Knowledge via Preference Optimization
In all of the aforementioned preference optimization methods, an implicit or explicit constraint exists to keep the model as close as possible to the reference model. Specifically, including KL-divergence regularization facilitates adjustments without deviating significantly from the reference model $\pi_{\text{ref}}$. To support this claim, Figure 5(a) shows the KL-divergence between the output distributions of the model and the reference model on the original dataset, measured as $\mathbb{E}_{x\sim\mathcal{D}_{\text{reg}}}\left[D_{\mathrm{KL}}(\pi_{\theta}(\cdot|x)||\pi_{\mathrm{ref}}(\cdot|x))\right]$, after fine-tuning for the typographic robustness task using different optimization methods.
Compared to Cross-Entropy optimization, our proposed methods result in lower divergence, demonstrating better retention of the model’s pre-trained knowledge. Figure 5(b) illustrates that we can control this deviation (i.e., retention of pre-trained knowledge) during fine-tuning by adjusting the $\beta$ hyperparameter, consistent with Eq. (3). All models in (a) and (b) were trained on the Imagenet100 and FOOD101 datasets, respectively, and the models in (b) were optimized using the IPO method.
(a) Our methods vs. Cross-Entropy optimization (CE).
(b) Effect of changing the hyperparameter $\beta$.
Figure 5: KL-divergence studies.
@@ 5 Conclusion
This paper introduces a novel approach for improving contrastive learning models through Preference Optimization (PO). Our approach systematically aligns model behavior with desired preferences, enhancing robustness against adversarial attacks and effectively mitigating biases in vision-language models like CLIP. Our experiments demonstrate improved performance over standard contrastive learning methods in adversarial robustness and fairness. This work could open new pathways for applying preference-based optimization to non-generative models.
Acknowledgments
This work was partially funded by the Federal Ministry for Digital and Transport (BMDV), Germany (“MoToRes”, 19F2271A) and by the Federal Ministry for Economic Affairs and Climate Action (BMWK), Germany (“ATTENTION!”, 01MJ22012D).
@@ Appendix A Theoretical Foundations of RLHF
RLHF extends the traditional reinforcement learning paradigm by incorporating human preferences into the learning process (Ouyang et al., 2022). This feedback can take the form of direct preference comparisons between different actions or sequences of actions, or it can involve explicit corrections when the agent’s behavior is undesirable. Human judgments are converted into a reward signal that helps the agent refine its policy. RLHF has gained attention in areas like natural language processing and fine-tuning large language models, where specifying a clear reward function is difficult and aligning the agent’s behavior with human values or intentions is critical. This method offers a more robust approach to training AI systems in tasks that require subjective judgment or nuanced decision-making.
@@
where $D_{\mathrm{KL}}(\pi_{\theta}^{\text{RL}}(y|x)\|\pi_{\theta}^{\text{SFT}}(y|x))$ represents the KL divergence between the RL policy and the SFT policy. The coefficient $\beta$ controls the trade-off between encouraging exploration and maintaining consistency with the supervised model. KL divergence serves two purposes: it acts as an entropy bonus, encourages exploration, and ensures that the RL policy stays within the supervised fine-tuned model.
@@ A.1 Deriving Preference Optimization as a Supervised Learning Framework
Direct Preference Optimization (DPO). DPO (Rafailov et al., 2024) leverages offline preference data to directly optimize the policy without relying on reinforcement learning-based methods like PPO. DPO demonstrates that the optimal solution to Eq. (20), $\pi_{\theta}^{*}$, satisfies the following equation:
@@
The preference probability $P_{\theta}$ is derived from pairwise comparisons using the Bradley-Terry model. Substituting Eq. (21) into Eq. (19) yields the following loss function for DPO:
@@
In Eq. (22), $\sigma$ denotes the sigmoid function, and $\mathcal{D}$ represents the dataset consisting of pairwise preferences. Each preference triplet $(x,y_{w},y_{l})$ includes a prompt $x$, a preferred response $y_{w}$, and a less preferred response $y_{l}$. This formulation allows DPO to optimize directly on pairwise preference data, avoiding the instability of RL methods while maintaining mathematical equivalence with traditional RLHF.
Identity Preference Optimization (IPO). As the original DPO loss function has shown limitations in practice, such as overfitting to the preference dataset (Gheshlaghi Azar et al., 2024), IPO was proposed as an extension. IPO introduces a regularization term to the DPO loss to mitigate overfitting by controlling the gap between the log-likelihood ratios of the preferred and dispreferred outputs for both the model and the reference model. The IPO loss function is defined as:
@@
Kahneman-Tversky Optimization (KTO). In language modeling, collecting large-scale preference datasets can be difficult, but it is relatively easier to determine whether an output is desired or undesired. This challenge has inspired KTO (Ethayarajh et al., 2024), a method that bypasses the need for detailed preference data by using a binary signal to guide optimization. KTO is based on prospect theory (Tversky, 2016), directly maximizing the utility of generations using a model of human decision-making.
@@
where $v_{\text{KTO}}(x,y;\beta)$ adjusts based on whether the output is desired or not:
@@
Here, $z_{\text{ref}}$ represents the reference reward, and $r_{\text{KTO}}(x,y)=\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}$. The weighting function $w(y)$ differentiates between desired and undesired outputs, where $\lambda_{D}$ and $\lambda_{U}$ are hyper-parameters controlling the weighting of each.
KTO’s key advantage is its ability to focus on maximizing the reward for desired outputs without inflating the KL divergence term, effectively balancing utility and regularization. This makes KTO a practical solution for real-world applications, where collecting detailed preference data is infeasible.